17. Taunton's is an all-equity firm that has 156,500 shares of stock outstanding. The CFO is considering borrowing $299,000 at 6 percent interest to repurchase 25,500 shares. Ignoring taxes, what is the value of the firm?
Answer #1

Share price = amount borrowed/shares repurchased = 299000/25500=11.72549

Value = share price*share number = 11.72549*156500=1835039.19